# Shapley Values-enabled Progressive Pseudo Bag Augmentation for Whole-Slide Image Classification

### v2

@@ 4.5.2 Sensitivity to Hyper-parameters
The optimal number of pseudo bags, $M_{max}$, varies across datasets due to differences in magnification levels of the patches and the sizes of tumor regions. As depicted in Fig. 5, on the CAMELYON-16 dataset, our model exhibits peak performance with an $M_{max}$ of approximately 9, experiencing a sharp decline in effectiveness when $M_{max}$ exceeds this value. Similarly, the ideal $M_{max}$ or the BRACS and TCGA-LUNG datasets is found to be 10 and 14, respectively. The smaller $M_{max}$ on CAMELYON-16 can be attributed to the prevalence of micro metastasis slides containing few positive instances, even at a 20× magnification. In such scenarios, pseudo bag augmentation must balance between introducing additional noise to the training set and enhancing training diversity. Conversely, the larger cancer (subtype) regions in the BRACS and TCGA-LUNG datasets allow for division into more pseudo bags without compromising stability.
@@ 4.5.3 Progressive Pseudo Bag Augmentation
To ascertain the efficacy of progressively increasing the pseudo bag count and refining the initial pseudo bag assignment, we carried out a series of experiments. For this, we set the pseudo bag increment $\Delta M$ to 4, with the training rounds, $n$ as 5 for the CAMELYON-16 and TCGA-LUNG datasets, and extended to 10 for the more intricate BRACS dataset. The results, as indicated in Table 5, show that models incorporating both progressive tactics achieve the highest levels of performance. The CAMELYON-16 dataset is particularly sensitive to the number of pseudo bags, necessitating precise calibration to prevent the introduction of undue noise. Conversely, the BRACS dataset’s sensitivity lies in the initial setup of pseudo bags, owing to the challenge of distinguishing between subtypes. A more sophisticated initial setup significantly aids the model in accurately recognizing positive instances, leading to enhanced performance.
@@ 5 Conclusion
In this study, we tackle attention-related challenges within multiple-instance learning for whole-slide image classification, particularly the extreme distribution of attention and misidentification of positive instances. To overcome these challenges, we introduce accelerated Shapley value, which quantifies the contribution of each instance, to estimate IIS for the first time. This novel approach facilitates a more logical allocation of pseudo bags. Furthermore, we present a progressive pseudo bag augmented multiple-instance learning framework that incorporates Shapley value-based IIS and utilizes the expectation-maximization algorithm. This approach systematically improves pseudo bag augmentation, thus significantly enhancing the efficacy of MIL. Extensive experiments on three publicly available datasets demonstrate that our methodology surpasses existing state-of-the-art techniques. Additionally, the Shapley value-based IIS offers valuable class-wise interpretability for pathological whole-slide images. In future research, we plan to investigate alternative metrics for the precise and efficient estimation of instance importance scores, aiming to further refine the robustness and versatility of our proposed framework.

### v4

@@ 4.5.2 Sensitivity to Hyper-parameters
The optimal number of pseudo bags, $M_{max}$, varies across datasets due to differences in magnification levels of the patches and the sizes of tumor regions. As depicted in Fig. 6, on the CAMELYON-16 dataset, our model exhibits peak ACC and AUC performance with an $M_{max}$ of approximately 8. Similarly, the ideal $M_{max}$ or the BRACS, TCGA-LUNG, and TCGA-BRCA datasets are found to be 10, 14, and 6, respectively. The smaller $M_{max}$ on CAMELYON-16 can be attributed to the prevalence of micro metastasis slides containing few positive instances, even at a 20$\times$ magnification. In such scenarios, pseudo bag augmentation must balance between introducing additional noise to the training set and enhancing training diversity. Conversely, the larger cancer (subtype) regions in the BRACS, TCGA-LUNG, and TCGA-BRCA datasets allow for division into more pseudo bags without compromising stability.
@@ 4.5.3 Progressive Pseudo Bag Augmentation
To ascertain the efficacy of progressively increasing the pseudo bag count and refining the initial pseudo bag assignment, we carried out a series of experiments. For this, we set the pseudo bag increment $\Delta M$ to 4, with the training rounds, $n$ as 5 for the CAMELYON-16, TCGA-LUNG, and TCGA-BRCA datasets, and extended to 10 for the more intricate BRACS dataset. The results, as indicated in Table 5, show that models incorporating both progressive tactics achieve the highest levels of performance. The CAMELYON-16 dataset is particularly sensitive to the number of pseudo bags, necessitating precise calibration to prevent the introduction of undue noise. Conversely, the BRACS dataset’s sensitivity lies in the initial setup of pseudo bags, owing to the challenge of distinguishing between subtypes. A more sophisticated initial setup significantly aids the model in accurately recognizing positive instances, leading to enhanced performance.
@@ 5 Conclusion
In this study, we tackle attention-related challenges within multiple-instance learning for whole-slide image classification, focusing on the issues of extreme attention distribution and misidentification of positive instances. To overcome these challenges, we introduce the accelerated Shapley value to quantify the contribution of each instance, pioneering the estimation of IIS for the first time. Compared with traditional pseudo bag assignment strategies like random splitting, our approach provides a more logical allocation of pseudo bags, effectively mitigating the mislabeling issue to some extent. Furthermore, we present a progressive pseudo bag augmented MIL framework that incorporates Shapley value-based IIS and utilizes the expectation-maximization algorithm. This framework systematically enhances pseudo bag augmentation, significantly boosting the efficacy of MIL. Extensive experiments on four publicly available datasets demonstrate that our methodology surpasses existing state-of-the-art techniques. Additionally, the Shapley value-based IIS offers valuable class-wise interpretability. Superior performance and interpretation help pathologists make more comprehensive and efficient diagnoses in clinical practice.
In future research, we aim to investigate alternative metrics for precise and efficient estimation of IIS, striving to enhance the robustness and versatility of our proposed framework.